OpenStudy (anonymous):

Solve: 3(x + 4) greater than or equal to x - 2 x (greater than or equal to) -7 x (greater than or equal to) 7 x (less than or equal to) -7 x (greater than or equal to) -3

OpenStudy (anonymous):

Is there a sign between 3(x+4) and x-2?

OpenStudy (anonymous):

oh yes sorry it is a - greater than or equal to sign-

OpenStudy (anonymous):

well first distribute the 3 to the x and 4

OpenStudy (anonymous):

So it would be 3x + 12 ≥ x-2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3(x+4) >= x-2 3x+4x >=x-2 7x>=x-2 7x-x >=-2 6x >=-2 x >=-2/6 x>=-1/3 it means x = -1/3 or x>-1/3

OpenStudy (anonymous):

now subtract x on both sides

OpenStudy (anonymous):

nevermind ezel2323 gave u the answer

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

wait but x > -1/3 isnt an answer

OpenStudy (anonymous):

ur welcom Andrewisbadatmath

OpenStudy (anonymous):

u have two value of x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

ur welcome